### v1

# A Knowledge Graph-based Retrieval-Augmented Generation Framework for Algorithm Selection in the Facility Layout Problem

Abstract
Selecting a solution algorithm for the Facility Layout Problem (FLP), an NP-hard optimization problem with a multiobjective trade-off, is a complex task that requires deep expert knowledge. The performance of a given algorithm depends on specific problem characteristics such as its scale, objectives, and constraints. This creates a need for a data-driven recommendation method to guide algorithm selection in automated design systems. This paper introduces a new recommendation method to make such expertise accessible, based on a Knowledge Graph-based Retrieval-Augmented Generation (KG-RAG) framework. To address this, a domain-specific knowledge graph is constructed from published literature. The method then employs a multi-faceted retrieval mechanism to gather relevant evidence from this knowledge graph using three distinct approaches, which include a precise graph-based search, flexible vector-based search, and high-level cluster-based search. The retrieved evidence is utilized by a Large Language Model (LLM) to generate algorithm recommendations with data-driven reasoning. The proposed KG-RAG method is compared against a commercial LLM chatbot with access to the knowledge base as a table, across a series of diverse, real-world FLP test cases. Based on recommendation accuracy and reasoning capability, the proposed method performed significantly better than the commercial LLM chatbot.
Keywords: Facility Layout Problem, Algorithm Selection, Knowledge Graph, Retrieval-Augmented Generation (RAG), Large Language Models (LLM), Recommender Systems.
1 Introduction
The Facility Layout Problem (FLP) is an important part of industrial engineering and operations research. [1] Because the operating cost and service performance are directly influenced by effective layout design. The suitable placement of facilities reduces the total operating costs by 20% to 50% [2]. The computational complexity of FLP stems from its combined discrete and continuous nature. In practice, this complexity is magnified by multi-objective trade-offs. For example, minimize material handling costs and maximize safety. This multi-objective problem is coupled with constraints such as aspect-ratio, adjacency and circulation constraints, and dynamic considerations such as reconfigurations to accommodate product-mix changes, demand changes, or equipment changes. These considerations render many FLP variants NP-hard, so exact mathematical programming is computationally tractable only for small or specially structured instances [3]. This FLP is also dynamic in nature because of evolving real-world demand. The problem gets further complicated because selecting a suitable optimisation algorithm for this complex problem is challenging and requires expert knowledge.
The Solution techniques have been evolving as the complexity of the FLP increases, as mentioned above. Initial solutions, such as quadratic assignment problem [4], mixed-integer programming and graph-theoretic formulations [5] for equal-area FLP. These are optimum guaranteed, but with exponential growth in computational time, limiting industrial applications to problems. As the problem size grows, researchers shifted to heuristic approaches, such as construction and improvement heuristics. Here, the construction heuristics build layouts from the ground up, for example, automated layout design program and computerized relationship layout planning [6] whereas the improvement heuristics build solutions incrementally, for example computerized relative allocation of facilities technique [6]. These techniques are likely to converge to suboptimal layouts. To enhance the ability of handling large and complex search spaces, various metaheuristic techniques, including particle swarm optimization, genetic algorithms, ant simulated annealing, colony optimization, and others, have been introduced [7]. These techniques increase the likelihood of avoiding local optima. Moreover, the hybrid metaheuristic approaches, such as biased random-key genetic algorithms with linear programming repair, have recently been developed as state-of-the-art solutions, with enhanced solution quality and scalability [8].
Despite these developments, the situation in FLP today is a broad and scattered collection of niche algorithms. This is typical of the no free lunch theorems, which state that when performance is averaged across all possible optimization problems, no single algorithm is universally superior. The theorems imply that an algorithm performs better on one class of problems and may perform weaker on others. For example, an algorithm that performs well in small-scale, static settings may be desired in large-scale, multi-objective, or dynamic conditions [9]. As a result, attention has turned to the meta-problem of algorithm selection, which involves selecting the algorithm, representation, and parameters best suited to an instance of information [21].
The early attempts to address this meta-problem were static decision-support methods. For instance, a GenOpt-based framework to guide practitioners in selecting from a wide range of general optimization algorithms. Their approach, which utilized flowcharts and choice matrices, was designed to assist novices by matching algorithm characteristics to problem types and performance requirements. While this represents a valuable step towards systematizing algorithm selection, the framework was not specific to the unique complexities of the FLP. Furthermore, like other static methods, it is not dynamic and cannot adapt to new algorithms or shifting problem definitions without manual updates [10]. To improve flexibility, research employed data-driven approaches, namely meta-learning, that predict algorithm performance from problem meta-features, for example, size, constraints, and statistical properties [11]. SATzilla is just an example, where it selects the fastest solver from a portfolio based on instance features and performs well in combinatorial optimisation [12]. This has evolved to AutoML, where algorithm selection is combined with hyperparameter tuning in the combined algorithm selection and hyperparameter framework [13]. While adaptive frameworks and AutoML have built significant capability, Large Language Models (LLMs) such as OpenAI’s GPT series and Google’s Gemini introduce new possibilities [14]. Large, general-purpose corpora-trained, these models are well-suited to acquiring language, reasoning, and in-context learning but fail on domain-specific problems such as FLP optimisation, hallucinations (providing plausible but incorrect responses) [15] and weak domain adaptation [16].
AI research has turned to grounding LLMs with structured, verifiable knowledge to address these weaknesses. Two key paradigms dominate Retrieval-Augmented Generation (RAG) and Knowledge Graphs (KG). RAG retrieves relevant, domain-specific content before generation, reducing hallucinations and enabling integration of up-to-date information without retraining [17]. KG organise domain entities and relationships, enabling explainable reasoning and mitigating data sparsity or cold-start issues. The relational pathways in KG enhance accuracy and interpretability, which are critical for decision support in technical domains [18]. A new approach integrates these paradigms. A KG-RAG [19] model uses a structured KG as the retrieval source for a RAG. The KG offers an explainable, linked domain model, and RAG allows LLMs to flexibly query and blend this knowledge into transparent, evidence-based proposals. This integration is suitable for comprehensive domain knowledge and justifiable output for complex recommendation tasks. The literature indicates a convergence of a long-standing operations research problem with a contemporary AI opportunity. FLP algorithmic fragmentation under the No Free Lunch principle renders selection a vital bottleneck. Current machine learning-based selectors, though adaptive, tend to operate as black boxes. Concurrently, the KG-RAG paradigm addresses LLM limitations of hallucinations and poor performance in domain adaptation while allowing for structured, transparent reasoning. KG-RAG is a promising basis for an intelligent, interpretable, and adaptive FLP algorithm recommendation method.
This paper introduces a dynamic and domain-specific solution using the KG-RAG framework to generate data-driven algorithm recommendations to solve the FLP, given a user-defined query and a domain-specific knowledge graph of historical solutions. Our approach begins with constructing a formal KG that captures the complex relationships between FLP instances, their unique characteristics, and the performance of various solution methodologies. To leverage this structured knowledge base, the method employs a multi-faceted retrieval pipeline that integrates precise graph queries, flexible vector search, and high-level cluster analysis to gather comprehensive evidence. An LLM then synthesizes this evidence to generate data-driven and explainable recommendations for a complete methodology, including the algorithm, algorithm parameters, a suitable problem representation, and a constraint handling technique. This work includes an intelligent feedback loop, which allows the method to function as a continuously learning environment by incorporating new user-submitted solutions, thus ensuring its long-term relevance and accuracy.
The rest of the paper has been structured as follows: Section 2 presents the proposed architecture and methodology, detailing the construction, domain-specific Knowledge Graph, and the multi-faceted retrieval and recommendation pipeline design. Section 3 provides a comparative experimental evaluation, benchmarking the KG-RAG method against a strong baseline LLM to empirically validate its performance on recommendation accuracy and its ability to generate data-driven, interpretable reasoning.
2 Architecture and methodology
The facility layout problem concerns placing a set of facilities (machines, departments, storage areas, etc.) inside a bounded site to optimise one or more performance measures. Practical FLP variations span multiple dimensions that materially affect solver choice and complexity. These begin with geometric representation, from the equal-area rectangles of early models [4] to the unequal area, aspect-ratio rectangles and general polygons of modern applications [1]. The problem scale is equally decisive [18], with the objective structure further constraining algorithm suitability, as single-objective methods, for example, minimizing material-handling cost, efficient area utilisation, minimizing material flow, may be ineffective for multi-objective trade-offs, for example, cost vs. safety vs. maintainability and vice versa [20]. Then the constraint coupling fundamentally shapes the feasibility landscape, including non-overlap, adjacency requirements, boundary constraints, and other geometric or operational limits. Together, these factors create a vast search space, leading to high variability in algorithm performance across different FLP instances. This diversity, as discussed in Section 2, underscores the necessity for a specialised algorithm recommendation method to adjust to each case’s unique features.
Figure 1: Example facility layout of production area [28]
Bridging this gap requires a systematic algorithm-selection approach that moves beyond generic recommendations. An ideal method must perform instance-aware matching, aligning suggestions with the specific scale, geometry, and constraints of the input account for objective type, ensuring alignment with single- or multi-objective goals and provide constraint-handling guidance, recommending methods such as repair operators, penalty functions, or exact feasibility checks. Beyond naming a method, it should supply validated algorithm parameters, such as population size for a genetic algorithm and starting temperature and cooling schedule for simulated annealing, as users often lack expertise in setting effective hyperparameters. For adoption, the method must ensure explainability and provenance, justifying recommendations with historical performance data, and deliver actionable outputs.
2.1 Architectural overview
A recommendation method is proposed to address the previously identified requirements. The method’s architecture, depicted in Figure 2, comprises several key components, each designed to address a specific research gap.
  • Initialisation of user query input: The process begins with the user defining their FLP. This component addresses the need for input flexibility, allowing users to specify diverse combinations of objectives, number of facilities, and constraints, moving beyond the rigid definitions.
  • Hybrid evidence retrieval: The method initiates a parallel, three-streamed retrieval process to overcome the limitations of static and single-faceted search methods. This component provides a comprehensive and adaptable search by collecting evidence from the Neo4j Knowledge Graph. This includes:
    • Graph-based search: A precise graph-based search using Cypher to find exact or structurally similar problems.
    • Vector-based search: A dynamic vector-based search to find conceptually related problems based on the user’s text description.
    • Cluster-based search: An aggregate analysis to identify high-level trends for the given problem type, for example, top algorithms for large-scale, multi-objective problems.
  • Evidence compilation: The results from all three retrieval channels are integrated and compiled into one contextualized dossier. The evidence is then organized into a close-fitting prompt template with the user’s original query to anchor the language model.
  • LLM-powered recommendation: To address the black box nature and the risk of LLM hallucinations, the evidence from all three retrieval channels is passed to this component. The results are compiled into a contextualized dossier that grounds the Google Gemini LLM. The LLM is prompted not to use its general knowledge but to act as an expert analyst, compiling the provided data into a final, formatted recommendation with a clear, evidence-driven explanation.
  • Continuous learning via feedback: The design has a feedback loop. When users add new solved problem cases, the method learns and incorporates this information smartly, enriching the KG. This makes the method continuously learn and improve, allowing it to make more accurate and suitable recommendations in the future.
Figure 2: High-level workflow of the recommendation
The method is implemented in Python, utilizing a modern technology stack composed of Neo4j for the graph database and vector index, Google Gemini as the LLM and embedding engine, LangChain for orchestrating the RAG pipeline, and Gradio for the interactive web user interface.
2.2 Knowledge graph construction
The method’s foundation is a designed KG that transforms a flat, tabular dataset into a rich, interconnected network of domain knowledge.
Table 1: Schema of a CSV source file
Column Type Example
problem_id string P_6, D_10
num_facilities integer 6, 10
floor_W, floor_H float 30, 90
Problem representation string continuous space
facility_dimension_data string fixed dim fixed area
objective string min material handling cost
constraints string non-overlapping
constraint_handling string shapely intersection
method string GA, PSO, TabuSearch
Model_parameters String pop_size=50;n_gen=200;
cost float 1147.781, 18446.919
time_sec float 185.0, 290.19
source(reference) String Knez, M. and Gajsek, B.,
2.2.1 Data corpus and schema
The initial knowledge base was derived from a CSV file, a comprehensive dataset aggregating solved FLP instances from academic literature and benchmarks. Each row represents a unique problem-solution pair, capturing critical attributes as detailed in Table 1.
2.2.2 Graph schema design
This tabular data, as Table 1 schema, is loaded into a Neo4j graph database through a custom Cypher script. The graph schema, used in Table 2, and  3 is defined to describe the domain entities and their complex relationships explicitly.
Table 2: Nodes and properties schema
Node label Key properties
Problem problem_id, num_facilities
Method name
Objective name
Constraint name
Representation name
ConstraintHandling name
Solution id (unique), cost, time_sec
ObjectiveCluster classification of objective
ScaleCluster Classification on problem’s size
ConstraintHandlingCluster category of cons_ handling
Table 3: Relationship schema
Relationship structure
(Solution) - [SOLVED] $\rightarrow$ (Problem)
(Solution) - [USED_METHOD] $\rightarrow$ (Method)
(Problem) - [HAS_OBJECTIVE] $\rightarrow$ (Objective)
(Problem) - [HAS_CONSTRAINT] $\rightarrow$ (Constraint)
(Problem) - [HAS_REPRESENTATION] $\rightarrow$ (Representation)
(Problem) - [CONS_HANDLING] $\rightarrow$ (ConsHandling)
(Problem) - [BELONGS_TO_SCALE] $\rightarrow$ (ScaleCluster)
(Method) - [IS_TYPE_OF] $\rightarrow$ (MethodCluster)
(ConsHandling) - [IS_TYPE_OF] $\rightarrow$ (ConsHandlingCluster)
(Problem) - [OBJECTIVE_CLUSTER] $\rightarrow$ (ObjectiveCluster)
2.2.3 Automated clustering and vector-based indexing
The KG is enriched with categorical clusters and vector-based embeddings for higher-level reasoning.
  • Automated clustering: Following the first import of data, a sequence of Cypher queries automatically clusters nodes according to deterministic rules. For instance, issues are clustered into a scale cluster by their number of facilities and into an objective cluster by whether the objective name contains a comma, which is a reasonable heuristic for multi-objective problems. This rule-based, automated mechanism brings consistency and scalability with the later additions of data appended.
  • Embedding generation: Each problem node is augmented with a vector embedding to enable conceptual search. Key attributes, such as objective and constraints, induce a full-text description property. The text is passed through the Google AI embedding model, and the resultant vector is cached in the node and indexed into Neo4j’s native vector index. The process is done once on new data, indexing all problems for similarity search.
2.3 The hybrid RAG recommendation pipeline
Whenever the user provides the query (number of facilities, objectives, constraints, and other specifications, if any), the method initiates a multi-stage pipeline to gather complete evidence before making a recommendation. After submission, the method normalizes and validates the inputs. The selections from the dropdowns are treated as validated entities. The free-text entries are treated as unvalidated entities. This distinction is critical as it dictates how the graph-based search query is constructed.
Adaptive query generation: The algorithm generates a set of evidence-gathering queries from the validated and unvalidated inputs. The initial step is exhaustive and serves as the basis for the rest of the steps of the RAG pipeline.Also, the population embedding process happens at method boot time and, significantly, is re-run for every new instance of the problem added by the user feedback loop.
2.3.1 Multi-faceted evidence retrieval
The method employs three parallel retrieval strategies to build a rich context for the LLM.
  • Graph-based search: The algorithm performs an exact, structured search in the KG. A Cypher query is dynamically built from user inputs. The query has a package of intelligent features loaded for exceptional cases.
    • Flexible filtering: The query applies an or constraint between constraints and goals to prevent the search from narrowing down too much. Besides that, it also dynamically regulates the quantity of facilities. Instead of searching for an exact amount, the Cypher query searches for problems in a ± 25% range of the number given by the user (i.e., between 0.75 * n and 1.25 * n). This dynamic range significantly increases the possibility of retrieving meaningful historical information, even if no exact number matches the facility number in the knowledge graph.
    • Performance-oriented ordering: The retrieved results are not random; a multi-factor scoring method ranks them to favour the most holistically suitable solutions. This relevance-first prioritises finding the correct solution over merely finding a high-performing one. The ORDER BY clause now prioritises results in a high-level order: first by the highest objective score (best matching the most user-specified objectives), second by the highest constraint score (best matching the most user-specified constraints). Only as a final resort does it fall back to using proximity to the user’s number of facilities, the minimum cost, and the minimum time as tiebreakers. This means that the highest-ranked results are high-performing and, most importantly, relevant to the user’s problem.
    • Multi-objective handling: If the user picks more than one objective, the query logic is then optimized to favour pulling solutions for issues that are tagged explicitly in the multi-objective category within the KG. This ensures that pulled examples are relevant to the problem of balancing competing objectives.
    • Adaptive fallback for unprecedented scale: The method fails gracefully when a user query is beyond the range of the current knowledge base. During boot time, the method preloads the maximum number of facilities in the KG. When its initial graph-based search fails and the requested number of facilities by the user is greater than this preloaded maximum, a fallback query is triggered programmatically. The query smartly broadens its search to find the most relevant large-scale precedents. It retrieves solutions from problems that meet two conditions. Problems with a facility count in the top quartile of the dataset’s known scale. Problems belonging to the top-level scale cluster.
  • Vector-based search: The user’s free-text problem description is translated into a vector representation. The vector then performs a similarity search against the indexed problem nodes in Neo4j. This returns conceptually similar problems.
  • Cluster-based search: To further support the provision of high-level statistical context, the method analyses the list of problem IDs found by the graph-based search (before applying the final limit). It does not aggregate queries to find the top three most common patterns to solutions for the Scale cluster and objective cluster associated with the user query. This gives strong trend-based evidence as well as the instance-level results.
2.3.2 LLM compilation and generation
All the top-ranked graph-based outcomes, top vector-based matches, and cluster analysis trends retrieved evidence are combined into one comprehensive context. The dossier is written in a strict prompt template, with the original user query submitted to the Google Gemini LLM. The prompt tells the model to write as a professional analyst, combining all the presented evidence to answer the form of two sections: a straightforward recommendation and a clear data-driven reasoning part, clearly explaining the recommendation using the retrieved data. Such an approach is required to guarantee the output’s trust and transparency. By tightly coupling the LLM with the provided context, the method architecturally rules out model hallucination risk and refrains from making recommendations on untested assumptions.
Finally, it incorporates a continuous learning loop to ensure the method remains current and avoids the static knowledge base problem. When users provide new solved problem instances, a backend process ensures stable and autonomous knowledge integration. This includes data standardization, such as combining multiple objectives into a single, canonical string to maintain uniformity in the KG. The method then utilizes Cypher’s MERGE command to intelligently update the graph, either linking to existing entities or creating new ones without duplication. This feature enables dynamic and continuous learning, enriching the knowledge base and allowing the method to make more accurate and relevant recommendations in the future.
3 Comparative evaluation
This section presents the experimental design and results of a comparative study conducted to evaluate the performance of our proposed KG-RAG method. The primary objective is empirically confirming that grounding an LLM in a well-structured, domain-specific Knowledge Graph leads to significantly accurate and interpretable recommendations than a baseline LLM approach using raw, structured data.
3.1 Evaluation methodology
The evaluation is designed as a direct comparative analysis between our proposed model and a strong baseline model, with five distinct test cases designed to cover a range of real-world FLP scenarios, from simple, single-objective problems to more complex multi-objective and ill-defined queries. Finally, Section 4 concludes by summarizing the paper’s primary contributions and key findings, discussing the method’s current limitations, and outlining promising directions for future research.
  • Baseline model: The baseline model consists of the same Google Gemini 1.5 Flash LLM as the proposed model, but instead of a KG, it is provided with the entire CSV dataset file as its knowledge source. This represents a powerful, traditional RAG method that relies on the LLM’s ability to find and reason with information from a flat, tabular text file.
  • Ground truth establishment: A ground truth recommendation was established for each test case to create a fair benchmark for accuracy. This was not a subjective choice but the result of a systematic, manual analysis of the source dataset, following a process that mirrors the logic of ground truth finding.
    Figure 3: Ground truth establishment flow
    Figure 4: User Interface
  • Metrics: Each method’s performance was measured against the ground truth using two key criteria.
    1. 1.
      Recommendation accuracy: binary measure (1 for a match, 0 for a mismatch) of whether the method’s top recommended algorithm(s) matched or included the ground truth.
    2. 2.
      Reasoning quality: A qualitative score from 1 (Poor) to 5 (Excellent) assessing the extent to which the LLM’s explanation was detailed, rich, and verifiably based on the evidence provided. The reasoning quality for both methods was evaluated by a separate instance of the Google Gemini 2.5 Flash LLM, which was prompted to act as an impartial judge and score the outputs based on the following criteria.
      1. (a)
        1 (Poor): Irrelevant or ambiguous explanation.
      2. (b)
        2 (Weak): Plausible but not specific.
      3. (c)
        3 (Acceptable): Reasonable but unsupported by cited evidence.
      4. (d)
        4 (Good): Applies to some numbers, but the analysis is superficial.
      5. (e)
        5 (Excellent): Combines several correct, cited evidence points to construct a firm, evidence-based conclusion.
3.2 Results and discussion
Table 4 compares outcomes of the proposed and baseline model with the ground truth and the data-driven ranking based on their explainability.
Table 4: Comparative evaluation of recommendation methods
Sl. Test case Ground truth Baseline Rec. Baseline Rating KG-RAG Rec. KG-RAG Rating
1 No of facilities = 10, Objective = min material handling cost, Constraints = non-overlapping, boundary constraints. CRO-SL ACO-FBS 4 CRO-SL, BRKGA 5
2 No of facilities = 15, Objective = max closeness rating, min material handling cost, Constraints = non-overlapping, boundary constraints, aspect ratio. HSA ACO-FBS 1 HSA, HGA 5
3 No of facilities = 30, Objective = min material handling cost, Constraints = non-overlapping, boundary constraint, area requirement, aspect ratio. BRKGA-LP, GA-LP BRKGA-LP 5 BRKGA-LP, GA-LP 5
4 No of facilities = 40, Objective = min material handling cost, Constraints = non-overlapping, boundary constraint, aspect ratio. BRKGA-LP ACO-FBS 3 BRKGA-LP, ACO-FBS. 5
5 No of facilities = 30, Objective = None, Constraints = None Construction Heuristic PROP1 4 BRKGA-LP, Construction Heuristic 5
3.2.1 Analysis of baseline method performance
While capable of reasoning, the baseline LLM’s recommendations achieve an overall accuracy of only 20% (1 out of 5). Its failures highlight the limitations of reasoning over structured tabular data. In Test cases 1 & 4, the baseline failed to recommend the ground truth algorithm because its reasoning was limited to a superficial raw text analysis. It could not perform the high-level frequency analysis (like the KG-RAG’s cluster analysis) needed to identify the most applied and successful methods. Moreover, it fails to interpolate the number of facilities when the exact number of facilities is not present in the dataset, as in test case 4. In Test case 2, the baseline’s recommendation of ACO-FBS for a multi-objective problem was factually incorrect. A query against the provided dataset would have revealed that ACO-FBS is only associated with Single-Objective problems. The baseline, lacking this explicit relationship, made a plausible but technically flawed suggestion, resulting in a reasoning score of 1.
Figure 5: Model recommendation rating analysis
3.2.2 Analysis of KG-RAG method performance
The KG-RAG method performed exceptionally well, achieving 100% recommendation accuracy and a perfect 5/5 reasoning score on all test cases. Its strength lies in its ability to synthesize multiple layers of evidence from the structured graph. In Test case 1, the KG-RAG model’s reasoning addressed specific problem instances and higher-level trends from the objective and scale clusters. This allowed it to correctly identify CRO-SL as the top contender, a conclusion the baseline missed. In Test case 2, the method’s ability to explicitly query for problems within the multi-objective cluster was critical. This ensured that the evidence provided to the LLM was highly relevant, leading to the correct recommendation of HSA. In Test case 5, even when the query did not contain an objective or constraints, the KG-RAG method provided a more nuanced and helpful recommendation. Analysing the broader context of 30-facility problems suggested both a simple heuristic (Construction Heuristic) and a more powerful metaheuristic (BRKGA-LP), giving the user a better understanding of the available trade-offs. This shows that the evaluation confirms that while a baseline LLM can reason from raw data to a limited extent, the KG-RAG model’s ability to leverage a structured, connected knowledge base allows it to perform an accurate analysis, yielding superior recommendations and data-driven explanations.
4 Conclusion
The paper introduced a new recommendation method based on a KG-RAG framework for FLP. This method successfully integrates the formalized reasoning strength of a Neo4j graph database with the advanced compilation strength of an LLM to generate explainable, high-quality recommendations.
The primary contribution of this paper is a multi-stage pipeline that integrates graph-based, vector-based, and cluster-based search approaches to construct a rich, evidence-supported context for every user query. The evaluation strongly confirmed the benefit of this approach. The proposed KG-RAG method significantly outperformed a baseline LLM with direct access to the raw data, achieving 100% recommendation accuracy against the ground truth, compared to 20% for the baseline. Furthermore, the quality of the data-driven reasoning was consistently rated as an average score of 5/5, far exceeding the baseline’s average score (3.4/5). This performance stems from the KG-RAG method’s ability to synthesize high-level trends from clusters with granular evidence from individual problem nodes, resulting in accurate, transparent, and reliable justifications. The risk of hallucination is substantially reduced by grounding the LLM in a domain-specific knowledge graph.
Despite these promising results, the proposed method has certain limitations. The method’s effectiveness fundamentally depends on the quality and scope of its underlying KG. It is a domain-specific tool for the FLP and recommendations for other domain optimization problems without a new, relevant knowledge base. Additionally, while the method includes a fallback mechanism for out-of-scope queries, for example, the problems with a number of facilities exceeding the KG’s included maximum data, its recommendations in these scenarios are based on broader categorical data rather than direct, comparable evidence, making them inherently less precise. Addressing this requires continuously expanding the knowledge base over time through the intelligent feedback loop. Furthermore, this paper prescribes a tangible and scalable way of building continuously updated methods. The collective feedback loop, which continuously feeds and aggregates new solutions populated by the users, maintains the knowledge graph updated over time. This makes the method a dynamic knowledge base that updates itself based on the user’s input. Future work will attempt to extend the framework to other challenging combinatorial optimization problems.
Acknowledgement
The authors express their sincere gratitude to TCS Research, Tata Consultancy Services Ltd., especially to the CTO team, for providing the internship opportunity and valuable guidance that greatly contributed to the completion of this work.